OpenStudy (anonymous):

Solve: 4(x + 6) = 6(x − 2)

OpenStudy (anonymous):

Expand the equation and collect the like terms together! Can you do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[4x + 24 = 6x - 12\] Now try solving it...get like terms on on one side and solve for x

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

is it 12?

OpenStudy (anonymous):

x=18

OpenStudy (anonymous):

\[4x - 6x = -12 -42\]\[-2x = -36\]\[x = \frac{ -36 }{ -2 }\]\[x= 18\]